# Identification of a Kalman Filter: Consistency of Local Solutions

### v1

@@ Abstract
Prediction error and maximum likelihood methods are powerful tools for identifying linear dynamical systems and, in particular, enable the joint estimation of model parameters and the Kalman filter used for state estimation. A key limitation, however, is that these methods require solving a generally non-convex optimization problem to global optimality. This paper analyzes the statistical behavior of local minimizers in the special case where only the Kalman gain is estimated. We prove that these local solutions are statistically consistent estimates of the true Kalman gain. This follows from asymptotic unimodularity: as the dataset grows, the objective function converges to a limit with a unique local (and therefore global) minimizer. We further provide guidelines for designing the optimization problem for Kalman filter tuning and discuss extensions to the joint estimation of additional linear parameters and noise covariances. Finally, the theoretical results are illustrated using three examples of increasing complexity. The main practical takeaway of this paper is that difficulties caused by local minimizers in system identification are, at least, not attributable to the tuning of the Kalman gain.
@@ 1 Introduction
Identifying a model from measurements is an important task, especially for designing model-based controllers. To efficiently apply such algorithms, three requirements are central: an accurate predictive model, efficient online state estimation, and, sometimes, uncertainty quantification. Regarding the first requirement, it is often a mixture of prior knowledge from physics-based modeling and data-driven modeling. A popular approach for this task is parametric system identification using Prediction Error Methods (PEM) (Ljung2002PEM) or Maximum Likelihood Estimation (MLE) (Astrom1979; Simpson2023). Regarding the second requirement, for Linear Time-Invariant (LTI) systems, online state estimation is often performed using Kalman filters (Anderson1979). Such a filter requires knowledge of the process and measurement noise covariance matrices, which are often difficult to derive from the system’s physics. Several approaches exist to estimate them from data (Abbeel2005; Odelson2006), but if they are entirely unknown, it is often preferable to estimate the Kalman gain directly with PEM or MLE, possibly jointly with other parameters (Kuntz2025). While these methods have strong statistical guarantees, they require solving a generally non-convex optimization problem to global optimality for these guarantees to hold. This is a limitation because derivative-based optimization algorithms can only guarantee convergence to a local minimizer. A natural question arises: can we still provide statistical guarantees for local minimizers?
@@
The asymptotic unimodularity of PEM and MLE has been proven for a few specific classes of Single-Input Single-Output (SISO) systems. A summary of classic results is given in Ljung1999SystemId. A notable one is asymptotic unimodularity for ARMA models (Astrom1974), which are black-box single-output autonomous LTI systems. Other results exist for SISO systems with specific structures or input design (Soderstrom1975; Goodwin2003; Zou2009; Eckhard2012). To the best of the authors’ knowledge, this paper provides the first asymptotic unimodularity result for multiple-output systems (apart from the trivial case of linear regression), and the first time-domain analysis of this problem.
@@
In Section 2, we define the stochastic system of interest and the associated identification problem. In Section 3, we formulate PEM for this problem, where only the Kalman gain is estimated. We also propose stability-enforcing constraints, and recall the state-of-the-art result on the consistency of the global solutions. In Section 4, we state the main results of this paper: asymptotic unimodularity and consistency of local minimizers. These theoretical results are illustrated with numerical examples in Section 5. Finally, in Section 6 we discuss possible extensions of these results to more general settings. We draw conclusions and discuss future research directions in Section 7.
@@ 3 The prediction error method
In our setting, the PEM (Ljung2002PEM) leads to the following optimization problem:
@@
From a more practical point of view, there exist several ways to impose such a uniform stability constraint. For example, in Kuntz2025, this is done via Linear Matrix Inequalities (LMI), and in Diehl2009a, the spectral radius is approximated by a smooth function that involves some Lyapunov equation. Similarly, we propose to ensure (5) as follows:
@@
for some choice of $\alpha>0$. Note that this formulation is equivalent to imposing $\rho_{\alpha}(A-LC)\leq 1$ where $\rho_{\alpha}(\cdot)$ is the smooth spectral radius approximation used in Diehl2009a. As remarked there, the constraint $P{\,\succcurlyeq\,}0$ is in fact never active (because $P{\,\succcurlyeq\,}I{\,\succ\,}0$ for any feasible point), so even though (6) is a nonlinear semi-definite program, it can be (almost) treated as an ordinary nonlinear program in practice.
@@
It is well known that the PEM is strongly consistent (Ljung1999SystemId, Theorem 8.2): the global minimizers of (3) converge almost surely (i.e., with probability one) to the true parameters $L^{\star}$ when $N$ goes to infinity. The proof relies on the fact that, almost surely, the objective function $V_{N}(L)$ converges uniformly to its expected value (Ljung1999SystemId, Lemma 8.2), and that $L^{\star}$ minimizes this expected value:
@@
Regarding the uniform convergence of $V_{N}(L)$, the proof relies on a lemma for stochastic dynamical systems presented in Ljung1999SystemId and repeated here in Appendix B.
@@ 4 Consistency of local solutions
In this section, we prove that strong consistency also holds for local minimizers of (3) that are in the interior of $\mathcal{L}$}. This relies on two main results: the first provides the limit of the objective function and its derivatives, and the second establishes the unimodularity of this limit.
@@
Theorem 5 (Unimodularity of the limit)
@@ 5 Numerical Examples
where $q(t)$ is the position of the particle at time $t$, $\mu>0$ is the friction coefficient, and $f_{k}\sim{\mathcal{N}}(0,\,\sigma_{f}^{2})$ is a random force that remains constant over each sampling period of length $\Delta t>0$. The measurements take the form $y_{k}=q(k\Delta t)+v_{k}$, where $v_{k}\sim{\mathcal{N}}(0,\,\sigma_{v}^{2})$ is the measurement noise. This system is discretized analytically into a discrete-time LTI model with $x_{k}=[q(k\Delta t),\dot{q}(k\Delta t)]^{\!\top}\!$, which we can put in the innovation form (1) after computing the true Kalman gain $L^{\star}$ from the Discrete Algebraic Riccati Equation (DARE) (Anderson1979). To estimate the gain, we solve the constrained PEM problem (6) with the Lyapunov-based stability constraint with some constant $\alpha\in(0,1)$. In practice, we optimize over $L$ only by eliminating $P$ via (6b). Derivatives are computed with CasADi, and an interior-point method (Nocedal2006) is implementedfootnotemark: with line-search and a Gauss-Newton Hessian approximation.
To explore the optimization landscape, we draw $50$ initial guesses uniformly in the feasible set ${\mathcal{L}}_{\alpha}$ and solve the problem for different data lengths $N$. Figure 2 shows the iterates and final solutions in the plane $(L_{11},L_{21})$ together with the stable region where $\rho(A-LC)<1$ and the associated subset where $\alpha{\mathrm{Tr}}{\left(P(L)-I\right)}\leq 1$. Note that the latter acts as a barrier for the former, which is consistent with the theory in Diehl2009a.
@@
Finally, as empirical evidence of the asymptotic unimodularity of the cost, we generate $50$ feasible initial guesses using Kalman filters with random matrices $Q$ and $R$ and observe that all initializations lead to the same solution (up to tolerance) when $N=100$.
@@ 6 Extensions
Regarding the joint estimation of $S$, the cost function in (3) must also be modified. Instead, the MLE approach for dynamical systems can be employed (Astrom1979; Simpson2023):
@@ Appendix B Uniform law of large numbers
Lemma 9 (Theorem 2B.1 in Ljung1999SystemId)

### v2

@@ Abstract
Prediction error and maximum likelihood methods are powerful tools for identifying linear dynamical systems and, in particular, enable the joint estimation of model parameters and the Kalman filter used for state estimation. A key limitation, however, is that these methods require solving a generally non-convex optimization problem to global optimality. This paper analyzes the statistical behavior of local minimizers in the special case where only the Kalman gain is estimated. We prove that these local solutions are statistically consistent estimates of the true Kalman gain. This follows from asymptotic unimodality: as the dataset grows, the objective function converges to a limit with a unique local (and therefore global) minimizer. We further provide guidelines for designing the optimization problem for Kalman filter tuning and discuss extensions to the joint estimation of additional linear parameters and noise covariances. Finally, the theoretical results are illustrated using three examples of increasing complexity. The main practical takeaway of this paper is that difficulties caused by local minimizers in system identification are, at least, not attributable to the tuning of the Kalman gain.
@@ 1 Introduction
Identifying a model from measurements is an important task, especially for designing model-based controllers. To efficiently apply such algorithms, three requirements are central: an accurate predictive model, efficient online state estimation, and, sometimes, uncertainty quantification. Regarding the first requirement, it is often a mixture of prior knowledge from physics-based modeling and data-driven modeling. A popular approach for this task is parametric system identification using Prediction Error Methods (PEM) (Ljung, 2002) or Maximum Likelihood Estimation (MLE) (Åström, 1979; Simpson et al., 2023). Regarding the second requirement, for Linear Time-Invariant (LTI) systems, online state estimation is often performed using Kalman filters (Anderson and Moore, 1979). Such a filter requires knowledge of the process and measurement noise covariance matrices, which are often difficult to derive from the system’s physics. Several approaches exist to estimate them from data (Abbeel et al., 2005; Odelson et al., 2006), but if they are entirely unknown, it is often preferable to estimate the Kalman gain directly with PEM or MLE, possibly jointly with other parameters (Kuntz and Rawlings, 2025). While these methods have strong statistical guarantees, they require solving a generally non-convex optimization problem to global optimality for these guarantees to hold. This is a limitation because derivative-based optimization algorithms can only guarantee convergence to a local minimizer. A natural question arises: can we still provide statistical guarantees for local minimizers?
@@
The asymptotic unimodality of PEM and MLE has been proven for a few specific classes of Single-Input Single-Output (SISO) systems. A summary of classic results is given in Ljung (1999, Section 10.5). A notable one is asymptotic unimodality for ARMA models (Åström and Söderström, 1974), which are black-box single-output autonomous LTI systems. Other results exist for SISO systems with specific structures or input design (Söderström, 1975; Goodwin et al., 2003; Zou and Heath, 2009; Eckhard et al., 2012). To the best of the authors’ knowledge, this paper provides the first asymptotic unimodality result for multiple-output systems (apart from the trivial case of linear regression), and the first time-domain analysis of this problem.
@@
In Section 2, we define the stochastic system of interest and the associated identification problem. In Section 3, we formulate PEM for this problem, where only the Kalman gain is estimated. We also propose stability-enforcing constraints, and recall the state-of-the-art result on the consistency of the global solutions. In Section 4, we state the main results of this paper: asymptotic unimodality and consistency of local minimizers. These theoretical results are illustrated with numerical examples in Section 5. Finally, in Section 6 we discuss possible extensions of these results to more general settings. We draw conclusions and discuss future research directions in Section 7.
@@ 3 The prediction error method
In our setting, the PEM (Ljung, 2002) leads to the following optimization problem:
@@
From a more practical point of view, there exist several ways to impose such a uniform stability constraint. For example, in Kuntz and Rawlings (2025), this is done via Linear Matrix Inequalities (LMI), and in Diehl et al. (2009), the spectral radius is approximated by a smooth function that involves some Lyapunov equation. Similarly, we propose to ensure (5) as follows:
@@
for some choice of $\alpha>0$. Note that this formulation is equivalent to imposing $\rho_{\alpha}(A-LC)\leq 1$ where $\rho_{\alpha}(\cdot)$ is the smooth spectral radius approximation used in Diehl et al. (2009). As remarked there, the constraint $P{\,\succcurlyeq\,}0$ is in fact never active (because $P{\,\succcurlyeq\,}I{\,\succ\,}0$ for any feasible point), so even though (6) is a nonlinear semi-definite program, it can be (almost) treated as an ordinary nonlinear program in practice.
@@
It is well known that the PEM is strongly consistent (Ljung, 1999, Theorem 8.2): the global minimizers of (3) converge almost surely (i.e., with probability one) to the true parameters $L^{\star}$ when $N$ goes to infinity. The proof relies on the fact that, almost surely, the objective function $V_{N}(L)$ converges uniformly to its expected value (Ljung, 1999, Lemma 8.2), and that $L^{\star}$ minimizes this expected value:
@@
Regarding the uniform convergence of $V_{N}(L)$, the proof relies on a lemma for stochastic dynamical systems presented in Ljung (1999, Theorem 2B.1) and repeated here in Appendix B.
@@ 4 Consistency of local solutions
In this section, we prove that strong consistency also holds for local minimizers of (3) that are in the interior of $\mathcal{L}$}. This relies on two main results: the first provides the limit of the objective function and its derivatives, and the second establishes the unimodality of this limit.
@@
Theorem 5 (Unimodality of the limit)
@@ 5 Numerical Examples
where $q(t)$ is the position of the particle at time $t$, $\mu>0$ is the friction coefficient, and $f_{k}\sim{\mathcal{N}}(0,\,\sigma_{f}^{2})$ is a random force that remains constant over each sampling period of length $\Delta t>0$. The measurements take the form $y_{k}=q(k\Delta t)+v_{k}$, where $v_{k}\sim{\mathcal{N}}(0,\,\sigma_{v}^{2})$ is the measurement noise. This system is discretized analytically into a discrete-time LTI model with $x_{k}=[q(k\Delta t),\dot{q}(k\Delta t)]^{\!\top}\!$, which we can put in the innovation form (1) after computing the true Kalman gain $L^{\star}$ from the Discrete Algebraic Riccati Equation (DARE) (Anderson and Moore, 1979). To estimate the gain, we solve the constrained PEM problem (6) with the Lyapunov-based stability constraint with some constant $\alpha\in(0,1)$. In practice, we optimize over $L$ only by eliminating $P$ via (6b). Derivatives are computed with CasADi, and an interior-point method (Nocedal and Wright, 2006) is implementedfootnotemark: with line-search and a Gauss-Newton Hessian approximation.
To explore the optimization landscape, we draw $50$ initial guesses uniformly in the feasible set ${\mathcal{L}}_{\alpha}$ and solve the problem for different data lengths $N$. Figure 2 shows the iterates and final solutions in the plane $(L_{11},L_{21})$ together with the stable region where $\rho(A-LC)<1$ and the associated subset where $\alpha{\mathrm{Tr}}{\left(P(L)-I\right)}\leq 1$. Note that the latter acts as a barrier for the former, which is consistent with the theory in Diehl et al. (2009).
@@
Finally, as empirical evidence of the asymptotic unimodality of the cost, we generate $50$ feasible initial guesses using Kalman filters with random matrices $Q$ and $R$ and observe that all initializations lead to the same solution (up to tolerance) when $N=100$.
@@ 6 Extensions
Regarding the joint estimation of $S$, the cost function in (3) must also be modified. Instead, the MLE approach for dynamical systems can be employed (Åström, 1979; Simpson et al., 2023):
@@ Appendix B Uniform law of large numbers
Lemma 9 (Theorem 2B.1 in Ljung (1999))